# C: A Coverage Path Planning Algorithm for Unknown Environments using Rapidly Covering Graphs

### v1

@@ 12                   Create a link node below $\hat{n}_{i}$ at a distance $w$;
$\bullet$ Set up TSP: Once $\mathcal{N}^{CH}_{i}$ is discovered, it is covered using a locally optimal TSP trajectory $\Lambda^{TSP}_{i}$. To compute $\Lambda^{TSP}_{i}$, the TSP is set up as follows. Since RCG is built with only a sparse set of nodes, it is important to create additional nodes in $\mathcal{N}^{CH}_{i}$ to compute $\Lambda^{TSP}_{i}$. Thus, nodes are added on each lap inside a coverage hole with the inter-node distance of $w$. Fig. 9(c) shows an example of node addition inside a coverage hole to compute $\Lambda^{TSP}_{i}$. These nodes are then connected to form a fully connected sub-graph $\overline{\mathcal{G}}^{CH}_{i}=(\overline{\mathcal{N}}^{CH}_{i},\overline{\mathcal{E}}^{CH}_{i})$. The node set $\overline{\mathcal{N}}^{CH}_{i}$ includes the current node $\hat{n}_{i}$, the goal node $\hat{n}_{i+1}$, the coverage hole nodes $\mathcal{N}^{CH}_{i}$ and the appended nodes. Each edge in ¯$\overline{\mathcal{E}}^{CH}_{i}$ is assigned the A [39] (shortest path) cost between its two end nodes.
1 Function ComputeTSPTrajectory
2 Inputs to iteration $i$: Current node $\hat{n}_{i}$; Goal node $\hat{n}_{i+1}$; $\mathcal{N}^{CH}_{i}$.
@@ 12       $n^{TSP}_{e}$ is not specified;
The start node of the TSP trajectory is set to $n^{TSP}_{s}=\hat{n}_{i}$ (Alg. 3 line 4), while its end node $n^{TSP}_{e}$ is determined as follows. If $\hat{n}_{i+1}$ is connected to an Open node which is not in the coverage hole then $n^{TSP}_{e}=\hat{n}_{i+1}$ (Alg. 3 lines 5-6). This allows for synergistic merging of $\Lambda^{TSP}_{i}$ to the back-and-forth trajectory through $\hat{n}_{i+1}$. The above also applies when $\hat{n}_{i+1}$ is adjacent to unknown area; i.e., $\mathcal{B}(\hat{n}_{i+1},w)\cap\mathcal{A}_{u}\neq\emptyset$. Else, if $\hat{n}_{i}$ is connected to an Open node which is not in the coverage hole then $n^{TSP}_{e}=\hat{n}_{i}$ (Alg. 3 lines 7-8). This allows for synergistic merging of $\Lambda^{TSP}_{i}$ to the back-and-forth trajectory through $\hat{n}_{i}$. Note that $\hat{n}_{i}$ cannot be adjacent to unknown area because it is visted by the robot. If none of the above conditions are true then $n^{TSP}_{e}$ is not specified.
$\bullet$ Solve TSP: Once the TSP is set up it is solved (Alg. 3 line 12) in an efficient manner to get the TSP trajectory $\Lambda^{TSP}_{i}$. For conditions $1$ and $3$ above (Alg. 3 lines 5-6 and 9-10, respectively), since $n^{TSP}_{e}\neq n^{TSP}_{s}$, a dummy node $\eta$ is added to convert the problem into a closed-loop TSP that starts and ends at $\eta$. The edge cost between $\eta$ and $\hat{n}_{i}$ is set as $u_{\eta,\hat{n}_{i}}=0$. For condition 1, the edge cost between $\eta$ and $\hat{n}_{i+1}$ is set as $u_{\eta,\hat{n}_{i+1}}=0$, while for condition 3, it is set as $u_{\eta,\hat{n}_{i+1}}=\infty$. The edge costs between $\eta$ and all the remaining nodes is set as $u_{\eta,n}=\infty,\forall n\in\overline{\mathcal{N}}^{CH}_{i}\setminus\{\hat{n}_{i},\hat{n}_{i+1}\}$. Note that $\eta$ is removed once the optimized node sequence $\Lambda^{TSP}_{i}$ is obtained. For condition 2 (Alg. 3 lines 7-8), since $n^{TSP}_{e}=n^{TSP}_{s}$, the TSP is closed-loop; thus no dummy node is needed.
A heuristic approach is used to solve for TSP as it is NP hard [51]. First, the nearest neighbor algorithm [51] is used to obtain an initial node sequence. Then, the 2-opt algorithm [51] is used to improve this initial solution and obtain the optimized node sequence. Fig. 9(d) shows an example of the TSP trajectory executed inside a coverage hole.
@@ IV Algorithm Analysis
[…]
@@
2Inputs to iteration $i$: starting location $p_{i-1}$; existing RCG $\mathcal{G}_{i-1}$; next waypoint $p_{i}$; goal node $\hat{n}_{i}$
3
4while $\mathcal{N}^{Op}_{i-1}$neq $\emptyset$ do
5       $\mathcal{A}_{d,i}\leftarrow\textbf{NavigateSense}(p_{i-1},p_{i})$; // navigates from $p_{i-1}$ to $p_{i}$ and discovers the environment
6       $\mathcal{F}_{i}\leftarrow\textbf{CreateSamplingFront}(\mathcal{A}_{d,i})$; // creates a sampling front upon reaching $p_{i}$
7       $\mathcal{S}_{i}\leftarrow\textbf{GenerateFrontierSamples}(\mathcal{F}_{i})$; // Generates frontier samples on each lap in $\mathcal{F}_{i}$
8       $\mathcal{G}^{\prime}_{i-1}\leftarrow\textbf{ExpandRCG}(\mathcal{G}_{i-1},\mathcal{S}_{i})$; // assigns $S_{i}$ as $\mathcal{N}_{new}$ and adds connecting edges to $\mathcal{G}_{i-1}$
9       $\mathcal{G}_{i}\leftarrow\textbf{PruneRCG}(\mathcal{G}^{\prime}_{i-1})$; // generates updated RCG by pruning inessential nodes and edges
10       $\big{(}\hat{n}_{i+1},p_{i+1}\big{)}\leftarrow\textbf{SelectGoalNode}(\mathcal{G}_{i},\hat{n}_{i})$; // selects goal node, if a dead-end exists then uses the dead-end escape function as in Alg 1
11       $\big{(}q_{i}(\hat{n}_{i}),\mathcal{N}^{Op}_{i},\mathcal{N}^{Cl}_{i}\big{)}\leftarrow\textbf{UpdateState}\big{(}q_{i}(\hat{n}_{i}),\mathcal{N}^{Op}_{i},\mathcal{N}^{Cl}_{i}\big{)}$; // updates the state of $\hat{n}_{i}$ as in Alg 2
12       $\mathcal{N}^{CH}_{i}\leftarrow\textbf{DetectCoverageHoles}(\mathcal{G}_{i},\hat{n}_{i},\hat{n}_{i+1})$;
13       if $\mathcal{N}^{CH}_{i}\neq\emptyset$ then
14             // coverage hole exists $\Lambda^{TSP}_{i}\leftarrow\textbf{ComputeTSPTrajectory}(\mathcal{N}^{CH}_{i},\hat{n}_{i},\hat{n}_{i+1})$;
15             $n^{TSP}_{e}\leftarrow\textbf{ExecuteTSPTrajectory}(\Lambda^{TSP}_{i})$;
16             $\hat{n}_{i}=n^{TSP}_{e}$; $\hat{n}_{i+1}=n^{TSP}_{e}$;
@@ 19      $i=i+1$;
To find any coverage holes around node $\hat{n}_{i}$, a recursive search is conducted for Open nodes in the neighborhood. This has the worst case complexity of $O(|\mathcal{N}_{i}|)$; however, in practice it is significantly smaller. If a coverage hole is detected, then a fully connected sub-graph $\overline{\mathcal{G}}^{CH}_{i}=(\overline{\mathcal{N}}^{CH}_{i},\overline{\mathcal{E}}^{CH}_{i})$ is formed in the coverage hole and a TSP is formulated. An initial solution of TSP is obtained using a nearest-neighbor based heuristic approach which has a complexity of $O\left(|\overline{\mathcal{N}}^{CH}_{i}|^{2}\right)$ [51]. Then, the 2-Opt greedy algorithm is applied to improve the solution which has a complexity of $O\left(|\overline{\mathcal{N}}^{CH}_{i}|\right)$ [51]. Typically, $|\overline{\mathcal{N}^{CH}_{i}}|\ll|\mathcal{N}_{i}|$, thus the complexity of coverage hole prevention is $O(|\mathcal{N}_{i}|)$. ∎
@@
The space complexity of the C algorithm is $O(|\mathcal{N}_{i}|)$.
@@
  • Finally, the coverage continues until all nodes of RCG are visited or a dead-end is reached. The robot escapes from a dead-end to the closest retreat node and resumes the back-and-forth coverage. If there is no retreat node available, it implies that the complete coverage is achieved.
The above procedures guarantee that C achieves complete coverage of the free space. ∎
@@
Figure 12: Performance comparison of C with the baseline algorithms.
@@ V-A Validation on a Simulation Platform
The performance of C is first evaluated by extensive high-fidelity simulations on a multitude of complex scenarios. A robotic simulation platform, called Player/Stage [52] is used for simulations, where Player is a robot server that contains models of different types of robots and sensors, while Stage is a highly configurable simulator of a variety of robotic operations. Several complex scenarios of dimensions $50m\times 50m$ are randomly generated with different obstacle layouts to test and validate the performance of C. These scenarios are considered as unknown to the robot at the beginning of the CPP operation. In order to detect and map obstacles during navigation, the autonomous robot is equipped with a laser sensor that has the FOV with a span of $360^{\circ}$ and a detection range of $15m$. The robot has motion constraints including the top speed of $0.5m/s$, max acceleration of $0.5m/s^{2}$, and the minimum turning radius of $0.04m$. The robot dynamically discovers the environment during navigation and generates adaptive coverage trajectory until complete coverage is achieved.
@@ V-A1 Baseline Algorithms
Seven baseline algorithms are selected for performance comparison including: PPCPP [13], $\varepsilon^{*}$ [3], BA [22], NNCPP [23], BM [24], BSA [25], and FSSTC [26]. Table I compares the key features of C with the baseline algorithms. All algorithms were deployed in C+++++ + and run on a computer with $2.60$ GHz processor and $32$ GB RAM. The sampling resolution for C and the grid resolution for the baseline algorithms are set as $1m$.
@@ V-A2 Performance Metrics
Figure 13: Coverage ratio versus noise.
Figure 14: Robot ROSMASTER X3 integrated with sensors.
@@ V-A3 Comparative Evaluation Results
For further validation, twenty diverse scenarios with complex obstacle layouts are randomly generated, as shown in Fig. 12(a). The performance metrics of all algorithms are recorded for these scenarios and plotted in Fig. 12(b). The results show a similar trend where C provides a superior solution quality in all metrics as compared to the baseline algorithms. The average computation time recorded in these scenarios at each iteration including the sampling, RCG construction, and waypoint generation is of the order of $10^{-2}s$. The average computation time to find a TSP trajectory for a coverage hole is $\sim 0.015s$.
@@ V-A4 Performance in the Presence of Uncertainties
The performance of C is further evaluated in presence of sensor uncertainties. These uncertainties can result in poor localization of the robot as well as misplacement of detected obstacles, which could impact the overall coverage performance. Therefore, to analyze the effects of sensor uncertainties, noise was injected into the measurements of range detector (laser sensor), the heading angle (compass), and the localization system of the robot. A laser sensor typically admits a measurement error of 1% of its operation range and a modestly priced compass can provide heading information as accurate as $0.5$ [53]. The above errors were simulated with an Additive White Gaussian Noise (AWGN) with standard deviations of $\sigma_{laser}=1.5cm$ and $\sigma_{compass}=0.5^{\circ}$, respectively. The accuracy of the robot localization system depends on many factors, such as applied techniques, hardware platforms and environmental conditions. For example, the accuracy of a GPS system can range from $0.02m$ to $0.12m$ [53]. Thus, the uncertainty due to localization system is studied using an AWGN with standard deviation ranging from $\sigma_{localization}=0.02m$ to $0.12m$.
@@
The above uncertainties have a direct effect on the coverage ratio, i.e., the percentage of area covered with respect to the actual area of the coverage space. To calculate the coverage ratio, the search area is partitioned into a grid with a resolution of $1m$, then the total number of obstacle-free cells covered by the robot are recorded and divided by the total number free cells. Fig. 13 shows the box plot of the coverage ratio for different noise levels over ten Monte Carlo runs for the two scenarios presented above. The horizontal line in a box is the median result of the Monte Carlo simulations, while the box shows the middle $50th$ percentile values of results.
V-B Validation by Real Experiments
The performance of C is also validated by real experiments in a $7m\times 7m$ lab space with several static obstacles. Fig 14 shows the robot, ROSMASTER X3 [54], used for coverage of this space. The robot is equipped with 1) a RPLIDAR S2L [54] with a range of $8m$ for obstacle detection, 2) MD520 motor with an encoder [54] for detection of wheel rotation and linear displacement, and 3) MPU9250 IMU [54] for detection of speed, acceleration, and orientation. The Gmapping algorithm [55] is used for indoor localization. The robot carries Jetson Nano minicomputer that collects sensor measurements and runs C for real-time control and navigation. Table II lists the sensing and computing systems on the real robot.
Fig. 15 shows the various snapshots of an experiment where the robot covers the lab space by running C onboard with adaptive coverage trajectory generation in real-time. Each snapshot consists of two images where i) the left image shows the actual robot covering the lab space and ii) the right image shows the discovered and unknown areas, RCG nodes statuses marked with different colors (Open nodes as green, retreat nodes as blue and goal node as pink), and the coverage trajectory executed by the robot. The buffer zones marked by orange color are added around the obstacles for safety of the robot considering inertia and localization error. Fig. 15(a) shows the start of the coverage process at the bottom left corner where the robot travels ahead of the starting point. Fig. 15(b) shows the instant when the robot reaches a dead-end surrounded by obstacle and covered area in the local neighborhood. In this case, it escapes from the dead-end by selecting the nearest retreat node as the next waypoint. Fig. 15(c) shows that the robot continues the back and forth coverage. Fig. 15(d) shows the instant when the robot detects a coverage hole and starts the TSP-based optimal coverage. Fig. 15(e) shows the TSP-based optimal coverage of another coverage hole. Finally, Fig. 15(f) shows that the robot achieves complete coverage, thus validating the effectiveness of C on real robots. The average computation time recorded in the real experiment at each iteration is of the order of $10^{-2}s$. The average computation time to find a TSP trajectory is $\sim 0.008s$.
V-C CPP Applications
V-C1 Coverage using Energy-Constrained Robots
In this application, the CPP problem considers robots with an energy constraint, i.e., limited battery capacity. Thus, the robot has to make multiple trips back to the charging station to recharge its battery. As such, the coverage is completed in multiple cycles, where each cycle consists of three trajectory segments: 1) advance, 2) coverage, and 3) retreat.
During coverage the robot continuously monitors the remaining energy so that it has sufficient energy left for returning to the charging station. If the energy level is low then it returns back for recharging. For simplicity the energy consumption of the robot is modeled as proportional to the trajectory length [41]. For C, the robot evaluates its energy content every time it reaches a node of RCG to make a decision to return back for recharging. Thus, while sitting at the current node $\hat{n}_{i}$, it computes the expected remaining energy upon reaching the next goal node $\hat{n}_{i+1}$. This is computed as the current remaining energy minus the expected energy needed to move to $\hat{n}_{i+1}$. Then it computes the retreat trajectory from $\hat{n}_{i+1}$ to the charging station using A [39]. If the expected remaining energy at $\hat{n}_{i+1}$ is less than the expected energy consumption of the retreat trajectory, then it retreats back to the charging station from $\hat{n}_{i}$, otherwise it moves to $\hat{n}_{i+1}$. Once the robot is recharged at the charging station, the robot selects the nearest open node on RCG as the start node of next round and advances to this node along the shortest A path.
@@ V-C1 Coverage using Energy-Constrained Robots
Range $8m$
Fig. 16 shows the comparative evaluation results of C with the baseline algorithms ($\varepsilon^{*}+$ [40] and energy constrained online path planning (ECOCPP) algorithm [42]) on a randomly generated $50\text{m}\times 50\text{m}$ complex scenario. The battery charging station is located at the bottom left corner. The robot is considered to have the total energy capacity of $360$ units. Figs. 16(a)-16(c) show the trajectories of C, $\varepsilon^{*}+$ and ECOCPP, respectively, at different cycles with the advance, coverage and retreat segments. Fig 16(d) shows the performance comparison of C with the baseline algorithms in terms of different performance metrics. It is seen that C provides clean coverage pattern from left to right, with minimum coverage time and overlappings, and leaving behind no coverage holes. Overall, C achieves the best performance in all metrics.

### v2

@@ 11        Create a link node below $\hat{n}_{i}$ at a distance $w$;
$\bullet$ Set up TSP: Once $\mathcal{N}^{CH}_{i}$ is discovered, it is covered using a locally optimal TSP trajectory $\Lambda^{TSP}_{i}$. To compute $\Lambda^{TSP}_{i}$, the TSP is set up as follows. Since RCG is built with sparse nodes, it is important to create additional nodes in $\mathcal{N}^{CH}_{i}$ to compute $\Lambda^{TSP}_{i}$. Thus, nodes are added on each lap inside a coverage hole with the inter-node distance equal to the sampling resolution $w$ (Alg. 3 line 2). Fig. 9(c) shows an example of node addition inside a coverage hole to compute $\Lambda^{TSP}_{i}$. These nodes are then connected to form a fully connected sub-graph $\overline{\mathcal{G}}^{CH}_{i}=(\overline{\mathcal{N}}^{CH}_{i},\overline{\mathcal{E}}^{CH}_{i})$. The node set $\overline{\mathcal{N}}^{CH}_{i}$ includes the current node $\hat{n}_{i}$, the goal node $\hat{n}_{i+1}$, the coverage hole nodes $\mathcal{N}^{CH}_{i}$ and the appended nodes. Each edge in $\overline{\mathcal{E}}^{CH}_{i}$ is assigned the A [hart1968formal] cost between its end nodes.
The start node of the TSP trajectory is set to $n^{TSP}_{s}=\hat{n}_{i}$ (Alg. 3 line 3), while its end node $n^{TSP}_{e}$ is determined as follows. Condition 1: If $\hat{n}_{i+1}$ is connected to an Open node which is not in the coverage hole then $n^{TSP}_{e}=\hat{n}_{i+1}$ (Alg. 3 lines 4-5). This allows for synergistic merging of $\Lambda^{TSP}_{i}$ to the back-and-forth trajectory through $\hat{n}_{i+1}$. The above also applies when $\hat{n}_{i+1}$ is adjacent to unknown area; i.e., $\mathcal{B}(\hat{n}_{i+1},w)\cap\mathcal{A}^{u}\neq\emptyset$. Condition 2: Else, if $\hat{n}_{i}$ is connected to an Open node which is not in the coverage hole then $n^{TSP}_{e}=\hat{n}_{i}$ (Alg. 3 lines 6-7). This allows for synergistic merging of $\Lambda^{TSP}_{i}$ to the back-and-forth trajectory through $\hat{n}_{i}$. Note that $\hat{n}_{i}$ cannot be adjacent to unknown area because it is visited by the robot. Condition 3: If none of the above conditions are true, then $n^{TSP}_{e}$ is not specified (Alg. 3 lines 8-9).
$\bullet$ Solve TSP: Once the TSP is set up it is solved (Alg. 3 line 11) in an efficient manner to get the TSP trajectory $\Lambda^{TSP}_{i}$. For conditions $1$ and $3$ above (Alg. 3 lines 4-5 and 8-9, respectively), since $n^{TSP}_{e}\neq n^{TSP}_{s}$, a dummy node $\eta$ is added to convert the problem into a closed-loop TSP that starts and ends at $\eta$. The edge cost between $\eta$ and $\hat{n}_{i}$ is set as $u_{\eta,\hat{n}_{i}}=0$. For condition 1, the edge cost between $\eta$ and $\hat{n}_{i+1}$ is set as $u_{\eta,\hat{n}_{i+1}}=0$, while for condition 3, it is set as $u_{\eta,\hat{n}_{i+1}}=\infty$. The edge costs between $\eta$ and all the remaining nodes is set as $u_{\eta,n}=\infty,\forall n\in\overline{\mathcal{N}}^{CH}_{i}\setminus\{\hat{n}_{i},\hat{n}_{i+1}\}$. Note that $\eta$ is removed once the optimized node sequence $\Lambda^{TSP}_{i}$ is obtained. For condition 2 (Alg. 3 lines 6-7), since $n^{TSP}_{e}=n^{TSP}_{s}$, the TSP is closed-loop; thus no dummy node is needed.
1Function ComputeTSPTrajectory
Input: Iteration $i$: Current node $\hat{n}_{i}$; Goal node $\hat{n}_{i+1}$; $\mathcal{N}^{CH}_{i}$.
@@ 11  $n^{TSP}_{e}$ is not specified; // condition 3
A heuristic approach is used to solve for TSP as it is NP hard [aarts2003]. First, the nearest neighbor algorithm [aarts2003] is used to obtain an initial node sequence. Then, the 2-opt algorithm [aarts2003] is used to improve this initial solution and obtain the optimized node sequence. Fig. 9(d) shows an example of the TSP trajectory executed inside a coverage hole.
@@ IV Algorithm Analysis
$|\mathcal{S}_{\mathcal{F}_{i}}|<\left\lfloor\frac{2r_{d}}{w}\right\rfloor\left\lfloor\frac{\ell_{i}^{max}}{w}\right\rfloor$, where $\lfloor\rfloor$ is the floor operation, $\left\lfloor\frac{2r_{d}}{w}\right\rfloor$ is the maximum number of laps in a sampling front, and $\left\lfloor\frac{\ell_{i}^{max}}{w}\right\rfloor$ is the maximum number of samples on the longest lap in $\mathcal{F}_{i}$ with length $\ell_{i}^{max}\in\mathbb{R}^{+}$. Note that since $\mathcal{F}_{i}$ is generated by sensing during navigation from $p_{i-1}$ to $p_{i}$, $\ell^{max}_{i}$ could be greater than $2r_{d}$.
@@
Input: Iteration $i$: starting location $p_{i-1}$; existing RCG $\mathcal{G}_{i-1}$; next waypoint $p_{i}$; goal node $\hat{n}_{i}$.
@@
2while $\mathcal{N}^{\text{Op}}_{i-1}\neq\emptyset$ do
  // Navigate from $p_{i-1}$ to $p_{i}$ and discover the environment
3  $\mathcal{A}_{d,i}\leftarrow\textbf{NavigateSense}(p_{i-1},p_{i})$;
4
5 if $Reached(p_{i})$ then
     // Create a sampling front at $p_{i}$
6     $\mathcal{F}_{i}\leftarrow\textbf{CreateSamplingFront}(\mathcal{A}_{d,i})$;
7
    // Generate frontier samples on laps of $\mathcal{F}_{i}$
8     $\mathcal{S}_{i}\leftarrow\textbf{GenerateFrontierSamples}(\mathcal{F}_{i})$;
9
    // Assign $\mathcal{S}_{i}$ as $\mathcal{N}_{\text{new}}$ and add connecting edges to $\mathcal{G}_{i-1}$
10     $\mathcal{G}^{\prime}_{i-1}\leftarrow\textbf{ExpandRCG}(\mathcal{G}_{i-1},\mathcal{S}_{i})$;
11
    // Prune inessential nodes and edges
12     $\mathcal{G}_{i}\leftarrow\textbf{PruneRCG}(\mathcal{G}^{\prime}_{i-1})$;
13
    // Select goal node or escape dead-end (Alg. 1)
14     $(\hat{n}_{i+1},p_{i+1})\leftarrow\textbf{SelectGoalNode}(\mathcal{G}_{i},\hat{n}_{i})$;
15
    // Update the state of $\hat{n}_{i}$ (Alg. 2)
16     $(q_{i}(\hat{n}_{i}),\mathcal{N}^{\text{Op}}_{i},\mathcal{N}^{\text{Cl}}_{i})\leftarrow\textbf{UpdateState}(q_{i}(\hat{n}_{i}),\mathcal{N}^{\text{Op}}_{i},\mathcal{N}^{\text{Cl}}_{i})$;
@@ 14     $(\hat{n}_{i+1},p_{i+1})\leftarrow\textbf{SelectGoalNode}(\mathcal{G}_{i},\hat{n}_{i})$;
    // Detect coverage holes
18     $\mathcal{N}^{\text{CH}}_{i}\leftarrow\textbf{DetectCoverageHoles}(\mathcal{G}_{i},\hat{n}_{i},\hat{n}_{i+1})$;
19
20    if $\mathcal{N}^{\text{CH}}_{i}\neq\emptyset$ then
        // Plan and execute TSP trajectory
21        $\Lambda^{\text{TSP}}_{i}\leftarrow\textbf{ComputeTSPTrajectory}(\mathcal{N}^{\text{CH}}_{i},\hat{n}_{i},\hat{n}_{i+1})$;
22        $n^{\text{TSP}}_{e}\leftarrow\textbf{ExecuteTSPTrajectory}(\Lambda^{\text{TSP}}_{i})$
23        $\hat{n}_{i}\leftarrow n^{\text{TSP}}_{e}$; $\hat{n}_{i+1}\leftarrow n^{\text{TSP}}_{e}$;
24
@@ 26    $i\leftarrow i+1$;
To find any coverage holes around node $\hat{n}_{i}$, a recursive search is conducted for Open nodes in the neighborhood. This has the worst case complexity of $O(|\mathcal{N}_{i}|)$; however, in practice it is significantly smaller. If a coverage hole is detected, then a fully connected sub-graph $\overline{\mathcal{G}}^{CH}_{i}=(\overline{\mathcal{N}}^{CH}_{i},\overline{\mathcal{E}}^{CH}_{i})$ is formed in the coverage hole and a TSP is formulated. An initial solution of TSP is obtained using a nearest-neighbor based heuristic approach which has a complexity of $O\left(|\overline{\mathcal{N}}^{CH}_{i}|^{2}\right)$ [aarts2003]. Then, the 2-Opt greedy algorithm is applied to improve the solution which has a complexity of $O\left(|\overline{\mathcal{N}}^{CH}_{i}|\right)$ [aarts2003]. Typically, $|\overline{\mathcal{N}^{CH}_{i}}|\ll|\mathcal{N}_{i}|$, thus the complexity of coverage hole prevention is $O(|\mathcal{N}_{i}|)$. ∎
@@
The space complexity of C is $O(|\mathcal{N}_{i}|)$.
@@
  • Finally, the coverage continues until all nodes of RCG are visited or a dead-end is reached. The robot escapes from a dead-end to the closest retreat node and resumes the back-and-forth coverage. If there is no retreat node available, it implies that complete coverage is achieved.
Note: While C provides complete coverage of the interior free regions of a map, it provides resolution completeness with respect to the RCG along obstacle boundaries. ∎
@@
Figure 12: Performance comparison of C with the baseline algorithms on different scenarios
Figure 13: Coverage ratio versus noise.
@@ V-A Validation on a Simulation Platform
The performance of C is first evaluated by extensive high-fidelity simulations on a multitude of complex scenarios. A robotic simulation platform, called Player/Stage [gerkey2003player] is used, where Player is a robot server that contains models of different types of robots and sensors, while Stage is a highly configurable simulator of a variety of robotic operations. Several complex scenarios of dimensions $50$ m $\times$ $50$ m are randomly generated with different obstacle layouts to test and validate the performance of C. These scenarios are considered as unknown to the robot at the beginning of the CPP operation. In order to detect and map obstacles during navigation, the autonomous robot is equipped with a laser sensor that has the FOV with a span of $360^{\circ}$ and a detection range of $15$ m. The robot has motion constraints including the top speed of $0.5$ m/s, max acceleration of $0.5$ m/s2, and the minimum turning radius of $0.04$ m. The robot dynamically discovers the environment during navigation and generates adaptive coverage trajectory until complete coverage is achieved.
@@ V-A1 Baseline Algorithms
Seven baseline algorithms are selected for performance comparison including: PPCPP [hassan2019ppcpp], $\varepsilon^{*}$ [song2018], BA [viet2013ba], NNCPP [luo2008bioinspired], BM [ferranti2007brick], BSA [gonzalez2005bsa], and FSSTC [gabriely2003competitive]. Table I compares the key features of C with the baseline algorithms. All algorithms were deployed in C++++ and run on a computer with $2.60$ GHz processor and $32$ GB RAM. The sampling resolution for C and the grid resolution for the baseline algorithms are set as $1$ m.
@@ V-A3 Comparative Evaluation Results
For further validation, a total of twenty four diverse scenarios with complex obstacle layouts are randomly generated, as shown in Fig. 12(a). The performance metrics of all algorithms are recorded for these scenarios and plotted in Fig. 12(b). The results show a similar trend where C provides a superior solution quality in all metrics as compared to the baseline algorithms. The average computation time recorded in these scenarios at each iteration including the sampling, RCG construction, and waypoint generation is of the order of $10^{-2}$ s. The average computation time to find a TSP trajectory for a coverage hole is $\sim 0.015$ s.
$\bullet$ Comparison with the Optimal TSP-Solution: To evaluate the quality of C paths, we computed the optimal coverage path with the shortest length for each of the above scenario by solving a TSP over the full environment map using the Concorde exact solver [David2005]. Figs. 10(b), 11(b) and 12(b) show the TSP path lengths and it is seen over all scenarios that C achieves path lengths close to the optimal TSP solution, while maintaining real-time applicability in unknown environments.
Figure 15: Robot ROSMASTER X3 integrated with sensors.
@@ V-A4 Performance in the Presence of Uncertainties
The performance of C is further evaluated in the presence of sensor uncertainties. These uncertainties can result in poor localization of the robot as well as misplacement of detected obstacles, which could impact the overall coverage performance. Therefore, to analyze the effects of sensor uncertainties, noise was injected into the measurements of range detector (laser sensor), the heading angle (compass), and the localization system of the robot. A laser sensor typically admits a measurement error of 1% of its operation range and a modestly priced compass can provide heading information as accurate as $0.5$ [paull2013auv]. The above errors were simulated with Additive White Gaussian Noise (AWGN) with standard deviations of $\sigma_{laser}=1.5$ cm and $\sigma_{compass}=0.5^{\circ}$, respectively. The accuracy of the robot localization system depends on many factors, such as applied techniques, hardware platforms and environmental conditions. For example, the accuracy of a GPS system can range from $0.02$ m to $0.12$ m [paull2013auv]. Thus, the uncertainty due to localization system is studied using an AWGN with standard deviation ranging from $\sigma_{localization}=0.02$ m to $0.12$ m.
@@
The above uncertainties have a direct effect on the coverage ratio, i.e., the percentage of area covered with respect to the actual area of the coverage space. To calculate the coverage ratio, the search area is partitioned into a grid with a resolution of $1$ m, then the total number of obstacle-free cells covered by the robot are recorded and divided by the total number free cells. Fig. 13 shows the box plot of the coverage ratio for different noise levels over ten Monte Carlo runs for the two scenarios presented above. The horizontal line in a box is the median result of the Monte Carlo simulations, while the box shows the middle $50^{th}$ percentile values of results.
@@
Range 88 m
V-B Validation by Real Experiments
The performance of C is also validated by real experiments in a $7$ m $\times$ 77 m lab space with several static obstacles. Fig 15 shows the robot, ROSMASTER X3 [robot], used for coverage of this space. The robot is equipped with 1) a RPLIDAR S2L [robot] with a range of $8$ m for obstacle detection, 2) MD520 motor with an encoder [robot] for detection of wheel rotation and linear displacement, and 3) MPU9250 IMU [robot] for detection of speed, acceleration, and orientation. The Gmapping algorithm [grisetti2007improved] is used for indoor localization. The robot carries Jetson Nano minicomputer that collects sensor measurements and runs C for real-time control and navigation. Table II lists the sensing and computing systems on the real robot.
Fig. 14 shows the various snapshots of an experiment where the robot covers the lab space by running C onboard with adaptive coverage trajectory generation in real-time. Each snapshot consists of two images where i) the left image shows the actual robot covering the lab space and ii) the right image shows the discovered and unknown areas, RCG nodes statuses marked with different colors (Open nodes as green, retreat nodes as blue and goal node as pink), and the coverage trajectory executed by the robot. The buffer zones marked by orange color are added around the obstacles for safety of the robot considering inertia and localization error. Fig. 14(a) shows the start of the coverage process at the bottom left corner where the robot travels ahead of the starting point. Fig. 14(b) shows the instant when the robot reaches a dead-end surrounded by obstacle and covered area in the local neighborhood. In this case, it escapes from the dead-end by selecting the nearest retreat node as the next waypoint. Fig. 14(c) shows that the robot continues the back-and-forth coverage. Fig. 14(d) shows the instant when the robot detects a coverage hole and starts the TSP-based optimal coverage. Fig. 14(e) shows the TSP-based optimal coverage of another coverage hole. Finally, Fig. 14(f) shows that the robot achieves complete coverage, thus validating the effectiveness of C on real robots. The average computation time in the real experiment at each iteration is of the order of $10^{-2}$ s. The average computation time to find a TSP trajectory is $\sim 0.008$ s.
V-C CPP Applications
V-C1 Coverage using Energy-Constrained Robots
In this application, the CPP problem considers robots with an energy constraint, i.e., limited battery capacity. Thus, the robot has to make multiple trips back to the charging station to recharge its battery. As such, the coverage is completed in multiple cycles where each cycle consists of three trajectory segments: 1) advance, 2) coverage, and 3) retreat.
During coverage the robot continuously monitors the remaining energy so that it has sufficient energy left for returning to the charging station. If the energy level is low then it returns back for recharging. For simplicity the energy consumption of the robot is modeled as proportional to the trajectory length [shnaps2016]. For C, the robot evaluates its energy content every time it reaches a node of RCG to make a decision to return back for recharging. Thus, while sitting at the current node $\hat{n}_{i}$, it computes the expected remaining energy upon reaching the next goal node $\hat{n}_{i+1}$. This is computed as the current remaining energy minus the expected energy needed to move to $\hat{n}_{i+1}$. Then it computes the retreat trajectory from $\hat{n}_{i+1}$ to the charging station using A [hart1968formal]. If the expected remaining energy at $\hat{n}_{i+1}$ is less than the expected energy consumption of the retreat trajectory, then it retreats back to the charging station from $\hat{n}_{i}$, otherwise it moves to $\hat{n}_{i+1}$. Once the robot is recharged at the charging station, the robot selects the nearest open node on RCG as the start node of next round and advances to this node along the shortest A path.
Fig. 16 shows the comparative evaluation results of C with the baseline algorithms ($\varepsilon^{*}+$ [shen2020] and energy constrained online path planning (ECOCPP) algorithm [dogru2022eco]) on a randomly generated $50$ m $\times$ $50$ m complex scenario. The battery charging station is located at the bottom left corner. The robot is considered to have the total energy capacity of $360$ units. Figs. 16(a)-16(c) show the trajectories of C, $\varepsilon^{*}+$ and ECOCPP, respectively, at different cycles with the advance, coverage and retreat segments. Fig 16(d) shows the performance comparison of C with the baseline algorithms in terms of different performance metrics. It is seen that C provides clean coverage pattern from left to right, with minimum coverage time and overlappings, and leaving behind no coverage holes. Overall, C achieves the best performance in all metrics.